OpenStudy (anonymous):

t=2r+3 5r-4t=6 I need to solve the system by substitution and check my answer, please help :(((

jimthompson5910 (jim_thompson5910):

5r-4t=6 5r-4(2r+3)=6 ... replace t with 2r+3 (since t = 2r+3) 5r - 8r - 12 = 6 solve for r and tell me what you get

OpenStudy (anonymous):

working on it now

jimthompson5910 (jim_thompson5910):

ok

OpenStudy (anonymous):

r=-6

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

now use that to find t

OpenStudy (anonymous):

now i replace it in the equation and solve for t ?

OpenStudy (anonymous):

t=-9?

jimthompson5910 (jim_thompson5910):

perfect r = -6 t = -9

OpenStudy (anonymous):

can you keep helping me with the rest ?

OpenStudy (anonymous):

r+s=-12 4r-6s=12

jimthompson5910 (jim_thompson5910):

how much more do you have?

OpenStudy (anonymous):

12 more

OpenStudy (anonymous):

as long as you can help me with like six

OpenStudy (anonymous):

just like you did on the first one

OpenStudy (anonymous):

please :((((((((((

jimthompson5910 (jim_thompson5910):

12 is quite a lot, but 6 sounds more manageable

OpenStudy (anonymous):

okay

jimthompson5910 (jim_thompson5910):

we'll see how it goes, but I'm not doing all 12

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

r+s=-12 4r-6s=12 solve by substitution

jimthompson5910 (jim_thompson5910):

solve the first equation for r r+s=-12 r = -12 - s then plug this into the second equation 4r-6s=12 4(-12 - s) - 6s = 12 -48 - 4s - 6s = 12 solve for s

OpenStudy (anonymous):

i did that one all by myself :) i didnt even look at the screen but now its elimination :(

OpenStudy (anonymous):

x+y=12 x-y=2 solve by elimination

jimthompson5910 (jim_thompson5910):

add the two equations to get x+y=12 x-y=2 -------- 2x+0y=14 so 2x = 14 -----> x = ??

jimthompson5910 (jim_thompson5910):

notice how the y terms cancel if and only if the y coefficients are equal and opposite

OpenStudy (anonymous):

7 :-))))))))))))

OpenStudy (anonymous):

YOU ARE AWESOME

jimthompson5910 (jim_thompson5910):

lol thx

jimthompson5910 (jim_thompson5910):

yes x = 7, then use this with any original equation to find y

OpenStudy (anonymous):

y=5 :)

jimthompson5910 (jim_thompson5910):

good

OpenStudy (anonymous):

i just did one by myself, im going to continue until i need you but dont forget about me

jimthompson5910 (jim_thompson5910):

ok that's great you're doing it on your own I'm sure you'll be fine

jimthompson5910 (jim_thompson5910):

and you can always ask questions here

OpenStudy (anonymous):

okay wait i did this one and w gave me -2 but i dont think thats correct

OpenStudy (anonymous):

2w+5y=-24 3w-5y=14 and im solving for elimination

jimthompson5910 (jim_thompson5910):

w = -2 is correct

jimthompson5910 (jim_thompson5910):

you got to trust yourself more

OpenStudy (anonymous):

i got rid of the 5 y's added the w's, it gave me 5w and then added the -24 +14 and it gave me -10

jimthompson5910 (jim_thompson5910):

yep 5w = -10 ----> w = -2

OpenStudy (anonymous):

ok thanks just making sure

OpenStudy (anonymous):

and y=5, correct?

OpenStudy (anonymous):

it doesnt fit in....

jimthompson5910 (jim_thompson5910):

no y = 5 is incorrect

OpenStudy (anonymous):

help please?

jimthompson5910 (jim_thompson5910):

2w+5y=-24 2(-2)+5y=-24 -4 + 5y = -24 5y = -24 + 4 5y = -20 y = -20/5 y = -4

jimthompson5910 (jim_thompson5910):

you could use the other equation too, either way you will get y = -4

OpenStudy (anonymous):

okay i need help with this one

OpenStudy (anonymous):

9a-3d=3 -3a+d=-1 solve by elimination

OpenStudy (anonymous):

do i have to multiply 3 to all of the parts in the second equation?

jimthompson5910 (jim_thompson5910):

yes you do, then add the equations

OpenStudy (anonymous):

cant be solved because ts dependent

jimthompson5910 (jim_thompson5910):

so are there no solutions? or infinitely many?

OpenStudy (anonymous):

no solutions....

OpenStudy (anonymous):

i mean it will always be 0

jimthompson5910 (jim_thompson5910):

if you multiply everything in equation 2 by 3, you get -9a+3d = -3 then add this to the first equation to get 9a-3d=3 -9a+3d=-3 ------------ 0a+0d=0 so we get 0 = 0

jimthompson5910 (jim_thompson5910):

0 = 0 is always true no matter what so that means you can pick any number you want for 'a' and you'll have a solution (note: d will depend on 'a') therefore there are infinitely many solutions because of that fact

OpenStudy (anonymous):

what does it mean when a system is inconsistant

jimthompson5910 (jim_thompson5910):

when there are no solutions (so the two lines are parallel and never intersect)

OpenStudy (anonymous):

6x-2y=11 -9+3y=16

jimthompson5910 (jim_thompson5910):

is there an x in the second equation?

OpenStudy (anonymous):

6x-2y=11 -9x+3y+16

OpenStudy (anonymous):

my bad

OpenStudy (anonymous):

..

jimthompson5910 (jim_thompson5910):

ok which method do you want to use

OpenStudy (anonymous):

elimination

OpenStudy (anonymous):

i think it is inconsistant cause i couldent get anything

jimthompson5910 (jim_thompson5910):

I don't think it's inconsistent

jimthompson5910 (jim_thompson5910):

hmm let me check

jimthompson5910 (jim_thompson5910):

ok yes it is inconsistent

OpenStudy (anonymous):

y=4-x 3x+y=6 solve by elimination

OpenStudy (anonymous):

im stuck, please help

jimthompson5910 (jim_thompson5910):

y=4-x y+x=4-x+x x+y = 4

jimthompson5910 (jim_thompson5910):

so we have this system x+y = 4 3x+y = 6

OpenStudy (anonymous):

then i multipied x times -3 and i was super confused

OpenStudy (anonymous):

help

jimthompson5910 (jim_thompson5910):

why not multiply either equation by -1 then add (this will eliminate the y terms)

OpenStudy (anonymous):

i got -1x-1y=-4 3x+y=-6

jimthompson5910 (jim_thompson5910):

now add the two equations

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

your my new favorite teacher

jimthompson5910 (jim_thompson5910):

thanks, I'm flattered